# Deep Context-Conditioned Anomaly Detection for Tabular Data

### v1

@@ 7.3. Quantifying the Value of Context
8. Conclusion
In this work, we introduced a context-conditional anomaly detection framework that addresses a fundamental limitation of traditional anomaly detection methods: their reliance on global decision boundaries that fail to capture contextual variations in heterogeneous data. We demonstrated that modeling conditional distributions $P(\mathbf{Y}\mid\mathbf{C})$ rather than joint distributions $P(\mathbf{Y},\mathbf{C})$ offers significant advantages. Our theoretical analysis, grounded in variance decomposition and principles from discriminative learning, showed that conditional modeling effectively isolates intra-context variability while reducing noise from inter-context variations, leading to more precise anomaly detection boundaries. The proposed CWAE model consistently outperformed both its non-contextual baseline (WAE) and SOTA methods on diverse benchmark datasets.
The implications of this work extend beyond the specific CWAE architecture to establish contextual learning as a valuable paradigm for anomaly detection in tabular data. Our findings demonstrate that many real-world datasets contain latent contextual structure that, when properly leveraged, can significantly enhance detection performance compared to traditional global approaches. The framework’s effectiveness across datasets with varying complexity levels highlights its broad applicability. Future research directions include developing more sophisticated context selection algorithms, exploring multi-context conditioning, and extending the framework to other data modalities.
9. Ethical Considerations
This research addresses several important ethical considerations related to anomaly detection systems and their deployment in real-world applications. Our work focuses on improving the accuracy and fairness of anomaly detection through contextual modeling, which has both positive implications and potential risks that must be carefully considered.
Privacy and Data Protection. The datasets used in this study include sensitive information such as demographic data (census), financial records (bank), and cybersecurity logs (beth, lanl). While we used publicly available datasets that have been appropriately anonymized, the contextual approach we propose could potentially be applied to more sensitive data in practice. Organizations implementing contextual anomaly detection must ensure compliance with privacy regulations such as GDPR and implement appropriate data protection measures, including differential privacy techniques where applicable.
Fairness and Bias Mitigation. A key ethical advantage of our contextual approach is its potential to reduce algorithmic bias by explicitly modeling different contexts rather than applying a single global decision boundary. Traditional anomaly detection systems may unfairly flag certain groups as anomalous due to differences in baseline behavior patterns. By conditioning on context features, our approach can learn appropriate norms for different user groups, potentially reducing false positives for underrepresented populations. However, careful selection of context features is crucial to avoid perpetuating existing biases or creating new forms of discrimination. We recommend thorough bias auditing when deploying contextual anomaly detection systems, particularly in high-stakes applications such as fraud detection or security screening.
Transparency and Explainability. While our contextual approach provides more nuanced anomaly detection, it also introduces additional complexity that may impact system interpretability. The use of context-specific thresholds and conditional modeling may make it more difficult for end users to understand why certain decisions were made. Organizations deploying such systems should invest in explainability tools and provide clear documentation of how contextual factors influence anomaly detection decisions, particularly in applications where individuals may be adversely affected by false positive classifications.

### v2

@@ 7.3 Quantifying the Value of Context
8 Conclusion
This work establishes contextual learning as a general paradigm for anomaly detection in tabular data. The proposed framework provides a unified probabilistic formulation, a principled context selection strategy, and a theoretical foundation for modeling conditional distributions $P(\mathbf{Y}\mid\mathbf{C})$. Through variance decomposition and discriminative learning principles, we show that conditioning effectively isolates intra-context variability while mitigating noise from inter-context differences, leading to more robust and precise anomaly detection.
We instantiate this paradigm through CWAE, a lightweight generative model that operationalizes the contextual learning framework for tabular domains. Across diverse datasets spanning finance, cybersecurity, demographics, and network intrusion detection, the CWAE instantiation consistently outperforms both its non-contextual baseline (WAE) and SOTA unconditional models, demonstrating the practical effectiveness and generality of contextual learning.
Future research will focus on extending contextual learning to multi-context conditioning, developing learnable context discovery mechanisms, and applying the framework to other data modalities such as text, time series, and multimodal environments.
Appendix A Appendix
This appendix provides comprehensive implementation and evaluation details to help facilitate the reproducibility of our experimental results.
A.1 Model Architectures
CWAE Architecture Details
The Conditional Wasserstein Autoencoder (CWAE) implements the following architecture:
Embedding Layers:
  • Type: Trainable embedding layers initialized randomly
  • Dimension: 16-dimensional embeddings for categorical features
  • Vocabulary Size: Determined by unique values in training data for each feature, plus 1 for handling unknown values at test time
  • Unknown Value Handling: Reserved index initialized to maximum values across embedding dimensions
Encoder:
  • Input Layer: Context embeddings ($\mathbb{R}^{c}$) concatenated with content embeddings ($\mathbb{R}^{d-c}$)
  • Hidden Layer 1: Linear transformation from (c+dc)(c+d-c) to 128 dimensions with ReLU activation
  • Hidden Layer 2: Linear transformation from 128 to 64 dimensions with ReLU activation
  • Latent Layer: Linear transformation from 64 to $z_{dim}$ dimensions, where $z_{dim}=64$ by default
  • Regularization: No bias terms used in linear layers
Decoder:
For each content feature $i\in\{1,\ldots,d-c\}$:
  • Input: Latent representation $z$ concatenated with context embeddings, dimension $\mathbb{R}^{z_{dim}+c}$
  • Hidden Layer 1: Linear transformation from $(z_{dim}+c)$ to 64 dimensions with ReLU activation
  • Hidden Layer 2: Linear transformation from 64 to 128 dimensions with ReLU activation
  • Output Layer: Linear transformation from 128 to vocabulary size for feature $i$, producing logits for cross-entropy loss computation
  • Regularization: No bias terms used in linear layers
CVAE Architecture Details
The Conditional Variational Autoencoder extends CWAE with probabilistic latent variables:
  • Encoder: Same architecture as CWAE encoder, but outputs both mean $\mu$ and log-variance $\log\sigma^{2}$ for the latent distribution
  • Latent Sampling: Reparameterization trick used: $z=\mu+\sigma\odot\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$
  • Decoder: Identical architecture to CWAE decoder
  • KL Divergence: Computed as $D_{KL}(q(z|x,c)||p(z))=-\frac{1}{2}\sum_{j=1}^{z_{dim}}(1+\log\sigma_{j}^{2}-\mu_{j}^{2}-\sigma_{j}^{2})$
CMDN Architecture Details
The Conditional Mixture Density Network models output distributions as Gaussian mixtures:
  • Encoder: Similar architecture to CWAE with context conditioning
  • Output: Predicts mixture parameters ($\pi_{k},\mu_{k},\sigma_{k}$) for $K=5$ mixture components
  • Mixture Weights: Normalized using softmax: $\pi_{k}=\frac{\exp(\alpha_{k})}{\sum_{j=1}^{K}\exp(\alpha_{j})}$
  • Loss Function: Negative log-likelihood of the Gaussian mixture: $\log\sum_{k=1}^{K}\pi_{k}\mathcal{N}(x|\mu_{k},\sigma_{k}^{2})$
A.2 Training Hyperparameters
Default Training Configuration
All models use consistent hyperparameters across experiments. This is shown in Table 6.
Parameter Value
Learning Rate 0.001
Optimizer Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$)
Batch Size 2048
Maximum Epochs 25
Early Stopping Threshold 0.0001 (on training loss)
Weight Decay 0 (no L2 regularization)
Gradient Clipping None
Random Seed 43
Table 6: Default training hyperparameters used across all experiments.
Model-Specific Adjustments
  • CWAE: MMD regularization weight $\lambda=1.0$
  • CVAE: KL divergence weight $\beta=1.0$ (standard VAE formulation)
  • CMDN: Early stopping threshold = -20.0 (adapted to negative log-likelihood scale)
  • CADI: Uses 100 isolation trees with automatic contamination parameter selection
A.3 Computational Environment
Hardware Specifications
Experiments were conducted on the following hardware:
  • GPU: NVIDIA A100 GPU (40GB memory) with CUDA 12.4 support
  • CPU: Multi-core processor for parallel data loading (4 workers)
  • RAM: Sufficient for batch size of 2048 across all datasets
  • Storage: SSD storage for efficient data I/O operations
Software Dependencies
Experiments were conducted with the following software dependencies. These are shown in Table 7
Package Version
Python 3.8+
PyTorch 2.6.0+cu124
CUDA 12.4
cuDNN 9.1.0.70
NumPy 2.0.2
Pandas 2.3.0
Scikit-learn 1.6.1
Matplotlib 3.9.4
Seaborn 0.13.2
Table 7: Software dependencies and versions used in experiments.
A.4 Joint Validation Loss Computation
Motivation
To enable fair comparison between joint models (learning $p(x,y)$) and conditional models (learning $p(x|y)$), we convert all models to evaluate on the joint probability $p(x,y)$, as these distributions have different scales.
Conversion Methodology
For contextual models learning $p(x|y)$, we apply the probability chain rule:
$p(x,y)=p(x|y)\cdot p(y)$ (2)
In log-space, this becomes:
$\log p(x,y)=-\log p(x|y)-\log p(y)$ (3)
Marginal Probability Estimation:
  1. 1.
    Estimate $p(y)$ empirically from training data: $\hat{p}(y=c)=\frac{\text{count}(c)}{N}$
  2. 2.
    Compute log probabilities: $\log\hat{p}(y=c)=\log(\text{count}(c))-\log(N)$
  3. 3.
    Handle unseen values: Use minimum observed probability as a fallback for rare categories
Sample-wise Correction:
For each validation sample $(x_{i},y_{i})$:
$\text{joint\_loss}_{i}=\text{model\_loss}_{i}+\left(-\sum_{j\in\text{context}}\log\hat{p}(y_{ij})\right)$ (4)
Batch-wise Average:
$\text{Joint Validation Loss}=\frac{1}{B}\sum_{i=1}^{B}\text{joint\_loss}_{i}$ (5)
where $B$ is the number of validation batches (limited to 10 batches for computational efficiency).
A.5 Datasets
A.5.1 Data Format
All datasets were preprocessed into a standardized format:
  • Anomaly Label: Binary column with values 0 (normal) and 1 (anomaly)
  • Feature Encoding: All features encoded as categorical integer indices
  • Missing Values: Handled via imputation or exclusion prior to experiments
A.5.2 Feature Encoding
  1. 1.
    Categorical Encoding: All features mapped to integer indices using label encoding
  2. 2.
    Index 0 Reserved: Designated for unknown or unseen categorical values during inference
  3. 3.
    Vocabulary Construction: Built exclusively from training data; validation and test sets may contain previously unseen values
  4. 4.
    No Numerical Scaling: Not required as features are processed through trainable embedding layers
A.5.3 Context/Content Feature Assignment
For each dataset and context configuration:
  1. 1.
    Context Features: Selected feature(s) used for conditional modeling
  2. 2.
    Content Features: Remaining features to be reconstructed by the model
  3. 3.
    Column Organization: Features organized as context features followed by content features
  4. 4.
    No-Context Baseline: All features treated as content (empty context set)
A.5.4 Train/Validation/Test Splits
All experiments use fixed pre-determined data splits:
  • Training Set: Contains only normal samples (anomaly=0)
  • Validation Set: Contains only normal samples, used for context selection
  • Test Set: Contains both normal and anomalous samples for evaluation
  • Split Ratios: Approximately 60% train, 20% validation, 20% test
  • Split Strategy: Random but fixed across all experiments to ensure fair comparison
A.5.5 Reproducibility Measures
  • Fixed Random Seed: Seed value of 43 used for all random operations
  • Data Loading: Training data shuffled; validation and test data processed in fixed order
  • Consistent Splits: Same train/validation/test partitions used across all model comparisons
A.6 Evaluation Metrics
A.6.1 Anomaly Scoring
For reconstruction-based models (CWAE, CVAE, CMDN), the anomaly score is computed as:
$\text{score}(x,c)=\sum_{i=1}^{d-k}\text{CE}(y_{i},\hat{y}_{i})$ (6)
where CE denotes cross-entropy loss, $y_{i}$ represents the true content feature value, $\hat{y}_{i}$ is the model’s reconstruction, and dkd-k is the number of content features.
A.6.2 Context-Specific Thresholding
  1. 1.
    Training Threshold: For each context value $c$, compute maximum training score: $H_{c}=\max_{(x,c)\in\text{train}}\text{score}(x,c)$
  2. 2.
    Normalization: Normalize test scores relative to training threshold: $R_{c}(x)=\frac{\text{score}(x,c)}{H_{c}}$
  3. 3.
    Global Threshold Search: Grid search over 100 uniformly-spaced thresholds in range $0,\max(R_{c})]$
  4. 4.
    AUC-ROC Computation: Use normalized ratios as continuous anomaly scores
A.6.3 Performance Metrics
  • AUC-ROC: Area under receiver operating characteristic curve
  • AUC-PR: Area under precision-recall curve
  • F1-Score: Harmonic mean of precision and recall at optimal threshold
  • Precision: $\frac{TP}{TP+FP}$ at optimal threshold
  • Recall: $\frac{TP}{TP+FN}$ at optimal threshold
Optimal threshold determined by maximizing F1-score on the test set.
A.7 Context Selection Procedure
A.7.1 Bilevel Optimization Framework
Outer Loop (Context Selection):
  • Candidate Set: All individual features tested as single-feature contexts
  • Baseline: No-context configuration included for comparison
  • Evaluation Protocol: Train each candidate for 1 epoch and compute joint validation loss
Inner Loop (Model Training):
  • Training Duration: Single epoch for context selection (early-epoch proxy)
  • Objective: Minimize training loss using standard hyperparameters
  • No Validation-Based Training: Validation set only used for evaluation, not for early stopping during context selection
Selection Criterion:
$c^{*}=\operatorname*{arg\,min}_{c\in\mathcal{C}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{val}}}[-\log p(x,y|\theta_{c})]$ (7)
where $\theta_{c}$ represents model parameters trained with context feature $c$, and $\mathcal{C}$ is the set of candidate context features.
A.7.2 Early Epoch Justification
Single-epoch evaluation is motivated by:
  • Computational Efficiency: Reduces context selection time by approximately 96% (25× speedup compared to full training)
  • Relative Ranking: Early training dynamics strongly correlate with final performance rankings (liu2019darts)
  • Empirical Validation: Figure 3 demonstrates that the early-epoch proxy achieves better final performance on 5 of 8 datasets
  • Established Practice: Similar approaches used successfully in neural architecture search (bender2018understanding), model pruning (you2020drawing), and noisy label learning (liu2020early)
A.8 Runtime Analysis
A.8.1 Training Time
Approximate training times per model/dataset combination (25 epochs on NVIDIA A100 GPU) shown in Table 8.
Dataset Training Time Samples Time/Epoch
bank 5 min 41,188 12 sec
beth 45 min 1,141,078 108 sec
census 12 min 299,285 29 sec
cmc 2 min 1,473 5 sec
kdd 40 min 1,014,535 96 sec
lanl 1.5 hrs 2,542,727 216 sec
sf 2 min 1,066 5 sec
spotify 8 min 113,550 19 sec
Table 8: Approximate training times for CWAE on each dataset.
A.8.2 Inference Time
Per-sample anomaly score computation:
  • Forward Pass: Less than 1 millisecond per sample with batch size 2048
  • GPU Acceleration: Significantly faster than CPU-based inference
  • Deterministic Scoring: CWAE requires single forward pass (no sampling) unlike stochastic models (CVAE)
A.8.3 Context Selection Time
For each dataset with $N$ candidate context features:
  • Single Context Training: 1 epoch $\approx$ 4% of full training time
  • Total Context Selection: $N\times$ (1-epoch time) + validation evaluation time
  • Example (census): 38 features × 29 sec/epoch $\approx$ 18.4 minutes
  • Validation Evaluation: Less than 1 minute per context (evaluated on 10 batches)